Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Box- 6 red and 2 white

Expert replies
by MURALIDHARGAJULA » Thu Jun 02, 2011 11:55 pm
A box contains 6 red and 2 white balls. Alice, Jane and Kate keep on drawing a ball in succession till a white ball is drawn. What is the probability that Jane draws white ball?

A) 3/14
B) 3/28
C) 15/28
D) 3/10
E) 9/28

OA E
Join the discussion
Source: — Problem Solving |

by cans » Fri Jun 03, 2011 12:13 am
A,J,K draw white ball in succession.
P(J draws)=p(RW) + p(rrrrw)= (6/8)*(2/7) + (6/8)(5/7)(4/6)(3/5)(2/4) = 3/14 + 3/28 = 9/28
IMO E
RW - Alice draws red, Jane draws white and game stops
rrrw, alice draws red, jane misses white and draws red, kate draws red, and den again alice draws red and this time jane draws white.
If my post helped you- let me know by pushing the thanks button ;)

Contact me about long distance tutoring!
[email protected]

Cans!!
Join the discussion

by Ozlemg » Fri Jun 03, 2011 12:51 am
Hi Cans,
Thank you for your explanation indeed.
I got the logic, no problem in there but I have a question :

P(J draws)=p(RW) + p(rrrrw)= (6/8)*(2/7) + (6/8)(5/7)(4/6)(3/5)(2/4) = 3/14 + 3/28 = 9/28


The fraction I colored red is the probability of drawing white out of 2 white balls. Isn't it?
It is not the descending probability of drawing red?
The more you suffer before the test, the less you will do so in the test! :)
Join the discussion

by cans » Fri Jun 03, 2011 3:24 am
Ozlemg wrote:Hi Cans,
Thank you for your explanation indeed.
I got the logic, no problem in there but I have a question :

P(J draws)=p(RW) + p(rrrrw)= (6/8)*(2/7) + (6/8)(5/7)(4/6)(3/5)(2/4) = 3/14 + 3/28 = 9/28


The fraction I colored red is the probability of drawing white out of 2 white balls. Isn't it?
It is not the descending probability of drawing red?
Yup it is probability of selecting a white ball out of remaining 4 balls :)
If my post helped you- let me know by pushing the thanks button ;)

Contact me about long distance tutoring!
[email protected]

Cans!!
Join the discussion

by vikram4689 » Fri Jun 03, 2011 7:38 am
Why the last case is not considered. According to me there are 3 cases

RW = 3/14
RRRRW = 3/28
RRRRRRRW = 1/28
Premise: If you like my post
Conclusion : Press the Thanks Button ;)
Join the discussion

by galaxian » Fri Jun 03, 2011 11:53 pm
We have only 6 Red Balls & hence no last case.
Join the discussion

by smackmartine » Sat Jun 04, 2011 12:10 am
cans wrote:A,J,K draw white ball in succession.
P(J draws)=p(RW) + p(rrrrw)= (6/8)*(2/7) + (6/8)(5/7)(4/6)(3/5)(2/4) = 3/14 + 3/28 = 9/28
IMO E
RW - Alice draws red, Jane draws white and game stops
rrrw, alice draws red, jane misses white and draws red, kate draws red, and den again alice draws red and this time jane draws white.
I am waiting to hear your quants score :)
Join the discussion

by cans » Sat Jun 04, 2011 8:49 am
smackmartine wrote:
cans wrote:A,J,K draw white ball in succession.
P(J draws)=p(RW) + p(rrrrw)= (6/8)*(2/7) + (6/8)(5/7)(4/6)(3/5)(2/4) = 3/14 + 3/28 = 9/28
IMO E
RW - Alice draws red, Jane draws white and game stops
rrrw, alice draws red, jane misses white and draws red, kate draws red, and den again alice draws red and this time jane draws white.
I am waiting to hear your quants score :)
Thanks smackmartine :) It feels so good to hear nice compliments like this :)
If my post helped you- let me know by pushing the thanks button ;)

Contact me about long distance tutoring!
[email protected]

Cans!!
Join the discussion

by sandeep800 » Sun Jun 05, 2011 10:45 am
U r great in Quant Cans :) Thanx
A body of clay, a mind full of play, a moment’s life - that is me!!!
Join the discussion